OpenStudy (anonymous):

I'm just so boggled. Can anyone help? Write the expression and solve: Three more than twice a number is at least 5.

OpenStudy (anonymous):

2x+3=5...would this be right?

OpenStudy (anonymous):

It says "at least five". That's the bit driving me bonkers.

OpenStudy (anonymous):

2x+3<or=5 making 2x=8 and x atleast 4 i believe

OpenStudy (anonymous):

You'd have to minus 3 from each side, so I think it's a > 1. Thank you so much!

OpenStudy (anonymous):

I read it this way: \[2x+3\geq{5}\]

OpenStudy (anonymous):

So \[\ge\] 1?

OpenStudy (anonymous):

\[x\geq{1}\] You can right click on my equation, hit show formating and then paste that into the window that pops up when you hit the "E Equation" button. Modify the code to see how your formated equation is displayed.

OpenStudy (anonymous):

Oooh, ok. Thank you so much, sir!